# Is General-Purpose AI Reasoning Sensitive to Data-Induced Cognitive Biases? Dynamic Benchmarking on Typical Software Engineering Dilemmas

### v1

@@ 7 Discussion
Implications for GPAI robustness. Three implications follow. First, correctness and robustness are separable: models can match Prolog on unbiased items yet flip under biased phrasing. Accuracy-only evaluations risk overstating reliability in user-facing contexts. Second, generic scaling (larger models or longer reasoning) is insufficient. Sensitivity increased with Prolog‐measured complexity for most biases, indicating that more compute does not uniformly suppress shortcut use. Third, mitigation should combine prompt and system interventions: i) bias‐aware instructions (e.g., “ignore recency, popularity, or outcomes; reason from rules”) and structured rationales that explicitly check for (common) biases; ii) cross-checks against a symbolic oracle (here, Prolog) before emitting a final choice; iii) self-consistency with debate or voting that penalizes bias-conditioned flips; and iv) training-time regularization that minimizes decision changes between paired biased/unbiased prompts.
Transferability of the Protocol. Nothing in our pipeline is tied to software engineering beyond the Prolog axioms and seed-prompt wording. Substituting domain rules (e.g., clinical guidelines, financial regulations) and domain-specific dilemmas yields an equivalent benchmark for medicine, law, or public policy. The Prolog round-trip test, similarity filters, and automatic bias adjudication are unchanged. Where Prolog is unsuitable, it can be replaced with a domain-appropriate executable oracle (e.g., a rules engine or SMT solver). When full formalization is impractical, partial specifications (e.g., constraints) still support neutrality and alignment checks, with weaker guarantees. Overall, the methodology is a general template for building dynamic bias-sensitivity benchmarks in any domain with a compact rule-based formalization.
Limitations. Our seed corpus and bias set are finite, and other corpora may yield different benchmarks. To mitigate this, we employ multiple data generators and confirm via diversity measures that coverage is good. The LLM-as-judge component is imperfect, especially when the same generator is used as judge, but also people are. Hence, we rely on a multi-step protocol that combines majority voting with manual validation of both dilemmas and Prolog programs, yielding high-quality items (92–99% correct). Additionally, Prolog complexity is only a proxy for cognitive load; however, our complexity analysis indicates it is a useful proxy. Finally, while Prolog captures rule-like reasoning, some practices may elude formalization; our program-dilemma alignment step is designed to exclude such cases.
8 Conclusion and Future Work
@@ Appendix A Bias Definitions
Table 1: Adopted definitions of cognitive biases
@@ Appendix B Hand-Crafted Seed Corpus: Extra Data
Figure 6: Hand-crafted seed corpus construction steps for biased/unbiased dilemmas, with bandwagon effect as the bias.
@@
For hyperbolic discounting, we used the inter-temporal choice framework typically employed in behavioural studies (Berns, Laibson, and Loewenstein 2007), where immediate smaller rewards are weighed against larger, delayed ones. Inspired by the methodology used by Benhabib, Bisin, and Schotter (2004) and case studies from P1, one dilemma (dubbed the bottleneck dilemma) depicted a situation where a developer’s quick fix result in code that hinder future progress. The unbiased version of this scenario was carefully adjusted to remove references to the speed or ease of the fix and any suggestion that the immediate solution might be preferable. Similarly, another dilemma was derived from a case where a developer initially opted for manual data collection due to unfamiliarity with a query interface (P1); the biased version emphasized the quickness of the manual approach (although it is not, in practice), while the unbiased version omitted such claims.
@@
Anchoring bias was examined through two scenarios that followed the anchoring-and-adjustment paradigm introduced by Tversky and Kahneman (1974). In one scenario, a suggested price or initial numerical estimate was provided, influencing subsequent valuation in a cost-versus-quality dilemma. In the other, a similar anchoring effect was observed in a time prediction scenario for a new implementation estimate, where the unbiased version is edited to remove any initial reference point.
Figure 7: Inter/intra-model agreement scores on seed corpus
@@
Finally, for the hindsight bias we adopted the retrospective judgment pattern (Chen et al. 2021): The corresponding dilemmas focused on scenarios such as new software product development and system refactoring, where a neutral (unbiased) or negative (biased) outcome was revealed after the initial decision. In the biased versions, the post-outcome information was framed to suggest significant failures (such as a major outage or intermittent system errors) even though the causal link to the developer’s actions was not explicitly confirmed, thus biasing the interpretation of the decision.
@@ 3 NL-to-Prolog Conversion
Convert a pair of natural-language dilemmas into Prolog code, separating the domain axioms from the specific facts of each variant.
@@
4 Prolog-to-Natural-Language Conversion
@@ 5 Bias-identification
6 Pairwise bias check
@@ Appendix D Data-Generator Selection Strategy
Appendix E Manual Benchmark Validation: Extra Data
Validators were given four CSV templates (task1_same_task.csv, task2_bias_presence.csv, task3_reconstruction.csv, and task4_axioms.csv) each containing items to be labeled True or False. They followed these criteria:
  • Axioms Description: Whether the best-practice axioms used in the Prolog program correctly conveys established software-engineering guidelines.
Agreement patterns by task and model. We compute percent agreement $P_{o}$ (Figure 8) and chance–corrected agreement (Fleiss’s $\kappa$) for each (generator, task) cell. Averaged across cells, $P_{o}=0.887$ and Fleiss’s $\kappa=0.194$. By task, agreement is highest for Bias Presence (mean $P_{o}=0.942$, range 0.9000.9000.9750.975) and lowest for Program-Dilemma Alignment (mean $P_{o}=0.808$, range 0.7500.7500.8500.850), with Same-Task Check and Axioms Appropriateness both at $P_{o}=0.900$ on average.
Figure 8: Raw percent agreement $P_{o}$ by generator and task.
Majority vs. unanimity pass rates. We summarize both majority-vote pass rate (threshold $\geq 2/3$ True) and a stricter unanimity metric (all three raters True). Averaged across generators and tasks, the majority-vote pass rate is $0.954$, while the unanimity pass rate is $0.877$. By task (averaging across generators), Same-Task Check and Bias Presence are near-ceiling under majority voting ($0.992$ and $0.983$, respectively) and remain high even under unanimity ($0.900$ and $0.942$). Program-Dilemma Alignment is the hardest: majority $0.883$, unanimity $0.767$. Axioms Appropriateness sits in between (majority $0.958$, unanimity $0.900$).
Model-wise trends. Averaged over tasks, gpt-4o-mini shows the highest majority pass rate ($0.969$) and unanimity rate ($0.906$), and also the highest mean $P_{o}$ ($0.906$) among generators. deepseek-r1 and gpt-4.1-mini are close (majority 0.9500.950/$0.944$; unanimity 0.8620.862/$0.862$; mean $P_{o}$ 0.8750.875/$0.881$). These gaps are visible in Figure 9 and align with the intuition that Program-Dilemma Alignment drives most of the between-model separation.
Figure 9: Comparison of majority-vote and unanimity pass rates by generator and task.
Takeaways for validation quality. (i) High majority-vote pass rates across Same-Task Check and Bias Presence confirm the robustness of those template families; (ii) Program-Dilemma Alignment is challenging; (iii) raw agreement is strong overall, while chance-corrected agreement is sensitive to skew, as expected with three raters; (iv) model trends are consistent across metrics, with gpt-4o-mini slightly ahead on both pass and agreement.
Appendix F Bias Sensitivity Analysis: Extra Data
Figure 10: Bias sensitivity analysis across common cognitive biases.
Appendix G Complexity Analysis: Extra Data
@@
Statistical Analysis. For each bias, we perform a one-sided two-proportion $z$-test with $H_{1}{:}\ p_{\text{high}}>p_{\text{low}}$, using aggregated binomial counts; 95% CIs for $p_{\text{high}}-p_{\text{low}}$ use the score (Newcombe) method (reported alongside one-sided $p$-values). As shown in Figure 11, the high complexity tier consistently exhibits greater sensitivity than the low tier across nearly all biases, with the estimated differences (in percentage points) lying well above zero and their 95% confidence intervals not overlapping the null. For most biases, the one-sided two-proportion $z$-test yields $p<0.05$, indicating a statistically significant improvement in sensitivity under increased complexity. A few biases display marginal effects (wider intervals and $p$-values near the threshold) reflecting smaller sample sizes or more variable performance. Overall, these results support the conclusion that higher dilemma complexity leads to higher bias sensitivity.
@@
Figure 11: High vs. low tier sensitivity difference with 95% CIs and one‐sided $p$-values.
@@
U-shaped complexity effect. In Figure 12, sensitivity often dips from mid-low to mid-high tiers before rebounding strongly at high complexity. All six models exhibit a positive swing of $\approx\!1\text{–}6\%$ at the highest tier relative to their own baselines, with GPT-4.1-nano showing the steepest jump (+5.6 pp).
@@
Figure 12: Aggregated bias sensitivity deviation from the average baseline sensitivity (unconditioned on complexity). The values shown are deltas from the overall mean sensitivity, i.e., the global average across all complexity tiers.
@@
Model-specific troughs. The heat-map (Figure 13) pinpoints where each model is most vulnerable: LLaMA-3.3 bottoms out at mid-high prompts (8.9%), whereas LLaMA-3.1 maintains the highest absolute sensitivity at both extremes (23.9% at low, 22.7% at high). GPT-4.1-mini remains the most stable, never exceeding 15%.
@@
Figure 13: Mean bias sensitivity (%) by model and complexity tier. Each cell shows the average sensitivity across all bias dimensions for a given model-tier pair.
@@
Heavy-tailed distributions. Figure 14 reveals widening, right-skewed tails as complexity rises, signalling that a handful of bias categories can elicit very large sensitivities (up to 100%). Conversely, the mid-low tier shows the tightest inter-quartile ranges, suggesting more predictable behaviour.
@@
Figure 14: Distribution of bias sensitivities (%) across individual bias dimensions, stratified by complexity tier and coloured by model.
@@
Practical implication. Safety evaluations that sample only “average-difficulty” prompts risk under-estimating real-world bias exposure. Including high-complexity scenarios is essential, and tuning efforts should prioritise the mid-low and high “valleys” where some models disproportionately under-perform.
Appendix H Protocol Scalability Analysis: Extra Data
@@
Table 2: Filter‐level failure counts and percentages by model.
@@
Appendix I Error Analysis of Bias‑Sensitive Entries
Figure 15: Heatmap of Pearson correlation coefficients between bias sensitivity and four proxy metrics across LLM variants.
[…]
Appendix J Thematic Analysis of Bias‑Sensitivity
To analyse how bias sensitivity manifests across biases, we conducted a qualitative analysis of dilemmas flagged as sensitive to bias. We employed an inductive (data-driven) thematic coding method, applied to decision justifications generated for both the biased and unbiased versions, to assess language-level differences. Due to the qualitative nature of this task, we analyzed only a subset of the dilemmas to extract emergent themes. Each theme was associated with one or more keywords manually identified in the samples. For instance, the theme security included terms like ’protection’, ’vulnerabilities’, and ’intrusion’. These keyword sets were then used to automatically label all bias-sensitive decision explanations.
@@ Appendix I Error Analysis of Bias‑Sensitive Entries
From this set, we inductively derived 14 high-level themes. Figure 16 shows the per-bias delta in theme counts (biased – unbiased) over the 40 samples. Figure 17 shows the absolute frequencies of themes across biased and unbiased conditions.
Figure 16: $\Delta$ Theme Counts (biased – unbiased) for each bias type on the 40 manually coded samples.
Figure 17: Absolute theme frequencies in biased vs. unbiased explanations (40 samples).
@@
Figure 18: $\Delta$ Theme Counts for All Models

### v2

@@ 10. Discussion
Implications for GPAI robustness.
Three implications follow. First, correctness and robustness are separable: models can match Prolog on unbiased items yet flip under biased phrasing. Accuracy-only evaluations risk overstating reliability in user-facing contexts. Second, generic scaling (larger models or longer reasoning) is insufficient. Sensitivity increased with Prolog‐measured complexity for most biases, indicating that more compute does not uniformly suppress shortcut use. Third, mitigation should combine prompt and system interventions: i) bias‐aware instructions (e.g., “ignore recency, popularity, or outcomes; reason from rules”) and structured rationales that explicitly check for (common) biases; ii) cross-checks against a symbolic oracle (here, Prolog) before emitting a final choice; iii) self-consistency with debate or voting that penalizes bias-conditioned flips; and iv) training-time regularization that minimizes decision changes between paired biased/unbiased prompts.
Transferability of the Protocol.
Nothing in our pipeline is tied to software engineering beyond the Prolog axiomatic background and seed-prompt wording. Substituting domain rules (e.g., clinical guidelines, financial regulations) and domain-specific dilemmas yields an equivalent benchmark for medicine, law, or public policy. The Prolog round-trip test, similarity filters, and automatic bias adjudication are unchanged. Where Prolog is unsuitable, it can be replaced with a domain-appropriate executable oracle (e.g., a rules engine or SMT solver). When full formalization is impractical, partial specifications (e.g., constraints) still support neutrality and alignment checks, with weaker guarantees. Overall, the methodology is a general template for building dynamic bias-sensitivity benchmarks in any domain with a compact rule-based formalization.
11. Conclusion and Future Work
@@ Appendix A Bias Definitions
Table 1. Adopted definitions of cognitive biases
@@ Appendix B Hand-Crafted Seed Corpus: Extra Data
Figure 9. Hand-crafted seed corpus construction steps for biased/unbiased dilemmas, with bandwagon effect as the bias.
@@
For hyperbolic discounting, we used the inter-temporal choice framework typically employed in behavioural studies (berns2007intertemporal), where immediate smaller rewards are weighed against larger, delayed ones. Inspired by the methodology used by benhabib2004hyperbolic and case studies from P1, one dilemma (dubbed the bottleneck dilemma) depicted a situation where a developer’s quick fix result in code that hinder future progress. The unbiased version of this scenario was carefully adjusted to remove references to the speed or ease of the fix and any suggestion that the immediate solution might be preferable. Similarly, another dilemma was derived from a case where a developer initially opted for manual data collection due to unfamiliarity with a query interface (P1); the biased version emphasized the quickness of the manual approach (although it is not, in practice), while the unbiased version omitted such claims.
@@
Anchoring bias was examined through two scenarios that followed the anchoring-and-adjustment paradigm introduced by tversky1974judgment. In one scenario, a suggested price or initial numerical estimate was provided, influencing subsequent valuation in a cost-versus-quality dilemma. In the other, a similar anchoring effect was observed in a time prediction scenario for a new implementation estimate, where the unbiased version is edited to remove any initial reference point.
Figure 10. Inter/intra-model agreement scores on seed corpus
@@
Finally, for the hindsight bias we adopted the retrospective judgment pattern (chen2021retrospective): The corresponding dilemmas focused on scenarios such as new software product development and system refactoring, where a neutral (unbiased) or negative (biased) outcome was revealed after the initial decision. In the biased versions, the post-outcome information was framed to suggest significant failures (such as a major outage or intermittent system errors) even though the causal link to the developer’s actions was not explicitly confirmed, thus biasing the interpretation of the decision.
@@ C.3. NL-to-Prolog Conversion
Convert a pair of natural-language dilemmas into Prolog code, separating the axiomatic background from the specific facts of each variant.
@@
C.4. Prolog-to-Natural-Language Conversion
@@ C.5. Bias-identification
C.6. Pairwise bias check
@@ Appendix D Data-Generator Selection Strategy
Appendix E Complexity Analysis: Extra Data
Statistical Analysis.
@@ Appendix E Complexity Analysis: Extra Data
For each bias, we perform a one-sided two-proportion $z$-test with $H_{1}{:}\ p_{\text{high}}>p_{\text{low}}$, using aggregated binomial counts; 95% CIs for $p_{\text{high}}-p_{\text{low}}$ use the score (Newcombe) method (reported alongside one-sided $p$-values). As shown in Figure 11, the high complexity tier consistently exhibits greater sensitivity than the low tier across nearly all biases, with the estimated differences (in percentage points) lying well above zero and their 95% confidence intervals not overlapping the null. For most biases, the one-sided two-proportion $z$-test yields $p<0.05$, indicating a statistically significant improvement in sensitivity under increased complexity. A few biases display marginal effects (wider intervals and $p$-values near the threshold) reflecting smaller sample sizes or more variable performance. Overall, these results support the conclusion that higher dilemma complexity leads to higher bias sensitivity.
@@
Figure 11. High vs. low tier sensitivity difference with 95% CIs and one‐sided $p$-values.
U-shaped complexity effect.
@@
In Figure 12, sensitivity often dips from mid-low to mid-high tiers before rebounding strongly at high complexity. All six models exhibit a positive swing of $\approx\!1\text{-}6\%$ at the highest tier relative to their own baselines, with GPT-4.1-nano showing the steepest jump (+5.6 pp).
@@
Figure 12. Aggregated bias sensitivity deviation from the average baseline sensitivity (unconditioned on complexity). The values shown are deltas from the overall mean sensitivity, i.e., the global average across all complexity tiers.
Model-specific troughs.
@@
The heat-map (Figure 13) pinpoints where each model is most vulnerable: LLaMA-3.3 bottoms out at mid-high prompts (8.9%), whereas LLaMA-3.1 maintains the highest absolute sensitivity at both extremes (23.9% at low, 22.7% at high). GPT-4.1-mini remains the most stable, never exceeding 15%.
@@
Figure 13. Mean bias sensitivity (%) by model and complexity tier. Each cell shows the average sensitivity across all bias dimensions for a given model-tier pair.
Heavy-tailed distributions.
@@
Figure 14 reveals widening, right-skewed tails as complexity rises, signalling that a handful of bias categories can elicit very large sensitivities (up to 100%). Conversely, the mid-low tier shows the tightest inter-quartile ranges, suggesting more predictable behaviour.
@@
Figure 14. Distribution of bias sensitivities (%) across individual bias dimensions, stratified by complexity tier and coloured by model.
Practical implication.
@@
Safety evaluations that sample only “average-difficulty” prompts risk under-estimating real-world bias exposure. Including high-complexity scenarios is essential, and tuning efforts should prioritise the mid-low and high “valleys” where some models disproportionately under-perform.
Appendix F Protocol Scalability Analysis: Extra Data
@@
Table 2. Filter‐level failure counts and percentages by model.
@@
Appendix G Thematic Analysis of Bias‑Sensitivity
To analyse how bias sensitivity manifests across biases, we conducted a qualitative analysis of dilemmas flagged as sensitive to bias. We employed an inductive (data-driven) thematic coding method, applied to decision justifications generated for both the biased and unbiased versions, to assess language-level differences. Due to the qualitative nature of this task, we analysed only a subset of the dilemmas to extract emergent themes. Each theme was associated with one or more keywords manually identified in the samples. For instance, the theme security included terms like ’protection’, ’vulnerabilities’, and ’intrusion’. These keyword sets were then used to automatically label all bias-sensitive decision explanations.
@@ Appendix G Thematic Analysis of Bias‑Sensitivity
From this set, we inductively derived 14 high-level themes. Figure 15 shows the per-bias delta in theme counts (biased - unbiased) over the 40 samples. Figure 16 shows the absolute frequencies of themes across biased and unbiased conditions.
Figure 15. $\Delta$ Theme Counts (biased - unbiased) for each bias type on the 40 manually coded samples.
Figure 16. Absolute theme frequencies in biased vs. unbiased explanations (40 samples).
@@
Figure 17. $\Delta$ Theme Counts for All Models